# Are Bus-Mounted Edge Servers Feasible?

### v2

@@ II Related Works
To achieve load balancing in the edge computing architecture of IoV, Hosono et al. proposed a collaborative load-balancing mechanism for edge servers based on Lane Section ID, which effectively mitigates local data processing overload caused by base station signal coverage deviation [8]. Feng et al. developed a joint offloading and resource allocation strategy for C-V2X vehicular networks, and designed a greedy offloading algorithm (PC5-GO) that dynamically coordinates the computation resources between RSUs and neighboring cooperative vehicles (CVs), achieving proactive load balancing and latency optimization under dual-interface constraints  [9].
@@ III-A Datasets
We use three datasets to evaluate the coverage and placement strategies of the buses: the Shanghai Telecom dataset [16, 17, 18], the Shanghai bus dataset [21], and the Shanghai taxi dataset [22]. The datasets span over a wide range of the Shanghai city, but we select the center of the map and partition it into grid cells with a side length of $200$ meters. The selected region is a square area that covers $72\times 72$ grids. The high density of the base stations, buses, and taxis in this area allows us to build a thorough understanding of the coverage of buses and base stations from the trace data.
@@
Shanghai Bus Dataset: This dataset includes over 55 million GPS records from over $2,100$ buses, during the period from February 19 to February 23 in 2007. In this dataset, we extract the longitude, latitude, time, and ID fields of the buses to construct their trajectories and derive the coverage. We treat these buses as potential edge servers which are able to handle computation requests generated by end users.
@@ III-B The Coverage of Potential Edge Servers
We divide the records in the trace into discrete time slots (e.g., 30 seconds each). For each time slot, we retrieve the locations of the base staions, buses, and taxis, and evaluate how many grids or taxis (demand points) can be covered by the base stations or buses, given a service radius $R$ of these potential servers. We introduce a family of metrics to quantify their coverage as follows.
@@
Point Coverage with Cap: This metric is similar to the previous metric. The difference is that Point Coverage with Cap enforces limited computation capacity of each server, while Point Coverage does not. If the number of demand points overwhelms the server capacity, they will not be covered even if they are within the service range of the server. In the right part of Fig. 1, the computation capacity of the server is 88 per time slot, so the 88 blue points are covered while the 55 black ones are not. This metric better reflects the actual ability of edge servers to meet the dynamic user demands.
@@ III-B1 Coverage of Base Stations
Fig. 2(a) illustrates the area coverage of the base stations, showing that the existing base station layout can cover most of the area. For instance, with a service radius of 11 km, $20\%$ of base stations yields nearly $90\%$ area coverage. Fig. 2(b) shows the point coverage, and the result also indicates a near-perfect coverage when the server computation capacity limit is ignored. However, Fig. 2(c) shows much worse results when we adopt a more realistic setting by limiting the server capacity to $50$ requests per time slot. The coverage decreases remarkably for all radii, which indicates that the capacity is in fact the primary bottleneck rather than the radius. It is worth noting that deploying edge servers with sufficient capacity to cover all demands all the time is irrational, as such capacity during normal periods may be wasted.
@@ III-B2 Coverage of Buses
The results show that buses also demonstrate desirable performance in terms of the three coverage metrics. The area and point coverage of buses, although smaller than base stations, can also reach over $50\%$ for certain server percentages and radii as shown in Fig. 3(a) and Fig. 3(b). Under more realistic scenarios shown in Fig. 3(c), the value of buses becomes pronounced. With a service radius greater than $0.8$ km and $50\%$ (approx. $1,070$) random buses equipped with edge servers, over $40\%$ of the demand points can be covered. Such coverage is closer to base stations compared with area coverage, taking into account that the total number of buses is only $2/3$ of the base stations. This suggests that buses have a great potential in acting as edge servers, compensating for the server capacity deficit of base station-based edge servers.
@@ V-B Performance Results
Fig. 6 shows a comparison of the demand volume covered by base stations and buses during the evening peak hour from 18:$0$ to 20:$0$, to show the effect of deploying both base station and bus-based edge servers. Fig. 6(a) shows the demand volume of the end users without any edge servers. Fig. 6(b) shows the remaining demand after deploying $30\%$ of the base stations, and Fig. 6(c) shows the result after introducing additional $500$ heuristically selected buses on that basis. Both the base stations and buses have a service radius of $0.8$ km and a capacity of $50$. The results reveal that while base stations handle most of the demands, some downtown areas still exhibit high load (the grids with red color), yet the buses effectively eliminate these remaining demand hotspots. This indicates that bus-mounted edge servers can cooperate with the base stations well, effectively handle the bursty traffic overwhelming the computation capacity of base stations during rush hours, and thereby improve the system computation elasticity and overall quality of service.

### v3

@@ II Related Works
Hosono et al. proposed a collaborative load-balancing mechanism for edge servers based on Lane Section ID, which effectively mitigates local data processing overload caused by base station signal coverage deviation [8]. Feng et al. developed a joint offloading and resource allocation strategy for C-V2X vehicular networks, and designed a greedy offloading algorithm (PC5-GO) that dynamically coordinates the computation resources between RSUs and neighboring cooperative vehicles (CVs), achieving proactive load balancing and latency optimization under dual-interface constraints  [9].
@@ III-A Datasets
We use three datasets to evaluate the coverage and placement strategies of the buses: the Shanghai Telecom dataset [16, 17, 18], the Shanghai bus dataset [21], and the Shanghai taxi dataset [22].
@@
Shanghai Bus Dataset: This dataset includes over $5$ million GPS records from over $2,100$ buses, during the period from February 19 to February 23 in 2007. In this dataset, we extract the longitude, latitude, time, and ID fields of the buses to construct their trajectories and derive the coverage. We treat these buses as potential edge servers which are able to handle computation requests generated by end users.
@@
The datasets span over a wide range of the Shanghai city, but we select the center of the map and partition it into grid cells with a side length of $200$ meters. The selected region is a square area that covers $72\times 72$ grids, containing $937$ base stations, $2,100$ buses, and $26$ million taxi GPS records. The high density of the base stations, buses, and taxis in this area allows us to build a thorough understanding of the coverage of buses and base stations from the trace data. We divide the records in the trace into discrete time slots (e.g., 30 seconds each). For each time slot, we retrieve the locations of the base staions, buses, and taxis, and evaluate how many grids or taxis (demand points) can be covered by the base stations or buses, given a service radius $R$ of these potential servers.
@@ III-B The Coverage of Potential Edge Servers
We introduce a family of metrics to quantify their coverage as follows.
@@
Point Coverage with Cap: This metric is similar to the previous metric. The difference is that Point Coverage with Cap enforces limited computation capacity of each server, while Point Coverage does not. If the number of demand points overwhelms the server capacity, they will not be covered even if they are within the service range of the server. In the right part of Fig. 1, the computation capacity of the server is $8$ per time slot, so the $8$ blue points are covered while the $5$ black ones are not. This metric better reflects the actual ability of edge servers to meet the dynamic user demands.
@@ III-B1 Coverage of Base Stations
Fig. 2(a) illustrates the area coverage of the base stations, showing that the existing base station layout can cover most of the area. For instance, with a service radius of $1$ km, $20\%$ of base stations yields nearly $90\%$ area coverage. Fig. 2(b) shows the point coverage, and the result also indicates a near-perfect coverage when the server computation capacity limit is ignored. However, Fig. 2(c) shows much worse results when we adopt a more realistic setting by limiting the server capacity to $50$ requests per time slot. The coverage decreases remarkably for all radii, which indicates that the capacity is in fact the primary bottleneck rather than the radius. It is worth noting that deploying edge servers with sufficient capacity to cover all demands all the time is irrational, as such capacity during normal periods may be wasted.
@@ III-B2 Coverage of Buses
The results show that buses also demonstrate desirable performance in terms of the three coverage metrics. The area and point coverage of buses, although smaller than base stations, can also reach over $50\%$ for certain server percentages and radii as shown in Fig. 3(a) and Fig. 3(b). Under more realistic scenarios shown in Fig. 3(c), the value of buses becomes pronounced. With a service radius greater than $0.8$ km and $50\%$ (approx. $1,070$) random buses equipped with edge servers, over $40\%$ of the demand points can be covered. Note that such point coverage ratio is based on the corresponding $27\%$-area coverage ratio. This indicates that the buses always go through high-traffic zones, which have a great potential in acting as edge servers and compensating for the server capacity deficit of base station-based edge servers.
@@ V-B Performance Results
Fig. 6 shows a comparison of the demand volume covered by base stations and buses during the evening peak hour from 18:00 to 20:00, to show the effect of deploying both base station and bus-based edge servers. Fig. 6(a) shows the demand volume of the end users without any edge servers. Fig. 6(b) shows the remaining demand after deploying $300$ random base stations, and Fig. 6(c) shows the result after introducing additional $200$ heuristically selected buses on that basis. Both the base stations and buses have a service radius of $0.8$ km and a capacity of $50$. The results reveal that while base stations handle most of the demands, some downtown areas still exhibit high load (the grids with red color), yet the buses effectively eliminate these remaining demand hotspots. This indicates that bus-mounted edge servers can cooperate with the base stations well, effectively handle the bursty traffic overwhelming the computation capacity of base stations during rush hours, and thereby improve the system computation elasticity and overall quality of service.